# How Far Can Transformers Reason? The Locality Barrier and Inductive Scratchpad

### v1

@@ 2.3 Agnostic scratchpads cannot break the locality
Consider training a $T$-regular net with an $s$-scratchpad to learn $P_{X,Y}$ on a constant-sized alphabet by means of the following SGD algorithm. At each timestep, we draw a random sample $(X,Y)$ and compute a value for the net with scratchpad on $X$. Let $\mathcal{S}$ be the resulting scratchpad and for each $i\leq s$ and each $\sigma\neq\mathcal{S}_{i}$, define an alternative scratchpad by setting the first $i-1$ entries of this scratchpad equal to those of $\mathcal{S}$, setting its $i$-th entry to $\sigma$, and using the net to compute the rest of its values. Then, regard the loss associated with setting the $i$th entry of the scratchpad to $\sigma$ as the loss resulting from the associated scratchpad, and use the resulting gradient to carry out one step of SGD. Then if $P_{X}$ is a well-behaved probability distribution, $P_{X,Y}$ is efficiently weakly learnable by a T-regular neural network with a scratchpad if and only if $P_{X,Y}$ has constant locality.
@@
3 Scratchpads to break the locality
@@ 3.1 Educated scratchpad
We now provide a quantitative understanding of how the scratchpad can help with the notion of locality in the autoregressive setting (Definition 3). Assume that we want to learn target $Y\in\mathcal{A}$ from input $X\in\mathcal{A}^{n}$ such that $(X,Y)\sim D$ and $\mathrm{Loc}(D)$ is high. If one can design intermediate targets $Y_{1},\ldots,Y_{k}\in\mathcal{A}$ such that $Y_{k}=Y$ and the sequence $(X,Y_{<i})\to Y_{i}$ has low locality according to Definition 3, then one can expect to learn each step of the sequence efficiently and thus the target at the end. In this case, the intermediate targets give the ‘educated scratchpad’ (see Figure 3 for an illustration). We now show how designing low-locality scratchpads can help with learning by focusing on two examples: parity functions and the cycle task.
Figure 3: An illustration showing how scratchpads can break the locality. The target may be efficiently learned if each scratchpad step is of low locality given the previous ones.
@@
Consider learning parity function $y=f(x_{1},\ldots,x_{n})=x_{1}x_{2}\cdots x_{k}$ where $x_{1},\ldots,x_{n}$ are drawn i.i.d. in {±$\{\pm 1\}$ with uniform distribution. For $k\leq\frac{n}{2}$, one can easily check that the locality of this task is $k$ as any $k-1$ coordinates are independent of the output and the histogram of the tokens does not help. Parity functions are known to be hard to learn [12]. More specifically, it has been previously shown that as $k$ increases the parity task becomes harder to learn to the point that parity of degree $k\wedge(n-k)=\omega(1)$ cannot be learned in $poly(n)$ time with standard poly-size neural networks under standard training assumptions [12]. Note that this is consistent with our results, as the locality is non-constant.
@@
The parity task with the cumulative product scratchpad has a locality of $2$.
@@
Consider the cycle task and a scratchpad that learns the depth-first search (DFS) algorithm from the source query node. [^9] For the particular graph structure of the two cycles task, DFS is the same as the breadth-first search (BFS). For example, consider the following input corresponding to two cycles $a,x,q$ and $n,y,t$: a>x; n>y; q>a; t>n; y>t; x>q; a?t;. In this case, doing a DFS from node a gives a>x>q>a where the fact that we have returned to the source node a and not seen the destination t indicates that the two nodes are not connected. Therefore, the full scratchpad with the final answer can be designed as a>x>q>a;0. Similarly, if the two nodes were connected the scratchpad would be a>...>t;1. One can easily check that the cycle task becomes low-locality with the DFS scratchpad.
@@
The cycle task with the DFS scratchpad has a locality of $3$.
@@
If one has full knowledge of the target function, one could break the target into sub-targets using an educated scratchpad to keep the locality low and thus learn more efficiently (of course one does not have to learn the target under full target knowledge, but one may still want to let a model learn it to develop useful representations in a broader/meta-learning context). One could in theory push this to learning any target that is poly-time computable by emulating a Turing machine in the steps of the scratchpad to keep the overall locality low. Some works have derived results in that direction, such as [33] for some type of linear autoregressive models, or [12] for more abstract neural nets that emulate any Turing machine with SGD training. However, these are mostly theory-oriented works. In practice, one may be instead interested in devising a more ‘generic’ scratchpad. In particular, a relevant feature in many reasoning tasks is the power of induction. This takes place for instance in the last 2 examples (parity and cycle task) where it appears useful to learn inductive steps when possible.
@@ 3.2 Inductive Scratchpads
As discussed previously, scratchpads can break the local reasoning barrier with appropriate mid-steps. In this part, however, we show that fully educated scratchpads can be sensitive to the number of reasoning steps, translating into poor out-of-distribution (OOD) generalization. As a remedy, we put forward the concept of inductive scratchpad which applies to a variety of reasoning tasks as in previous sections.
@@ 3.2.2 Inductive scratchpad: definition and experimental results
We can use inductive scratchpads to achieve length generalization for the parity and addition tasks. For parities, we insert random spaces between the bits and design an inductive scratchpad based on the position of the bits and then compute the parity iteratively. Using this scratchpad we can train a Transformer on up to $30$ bits and get generalization on up to $50$, or $55$ bits. Results are provided in Figure 5(a). For the addition task, we consider two inductive scratchpad formats. We provide an inductive scratchpad that requires random spaces between the digits in the input and uses the position of the digits to compute the addition digit-by-digit (similar to the parity). With this scratchpad, we can generalize to numbers with 20 digits while training on numbers with up to 10 digits. In the other format, we use random tokens in the input and we compute the addition digit-by-digit by shifting the operands. The latter enables us to generalize from 4 to 26 digits at the cost of having a less natural input format. The results for different seeds are provided in Figure 5(b). See details of these scratchpads in Appendices B.4B.5.101010Our code is also available at https://github.com/aryol/inductive-scratchpad. Also see Appendix A for a comparison with recent methods.
@@ 4 Conclusion
This paper shows that for the learning objective and in contrast to expressivity results, Transformers trained from scratch have a ‘local reasoning barrier’ quantified by the distribution locality: if the target requires a global composition of the inputs, Transformers will not efficiently learn. The locality measure has a simpler form and broader applicability range than prior measures as discussed in Appendix A.2, it also has tighter applicability for Transformer. The measure is currently defined for weak learning (inverse-polynomial or constant edge), and a natural next step is to consider stronger learning requirements with ‘locality leap’, e.g., expanding the current work in the direction of [28]. Investigating the role of curriculum learning is another natural direction and we provide preliminary results here in Appendix B.2.
The locality is also defined in the autoregressive setting, to better quantify when scratchpads can break targets into easier sub-targets. Two negative results are shown for scratchpads: agnostic scratchpads still suffer from the locality barrier, and fully educated scratchpads can have poor OOD generalization. This motivates the introduction of the inductive scratchpad.
The inductive scratchpad can be used for a broad range of reasoning/algorithmic tasks and can easily be integrated into Transformers. The inductive scratchpad is independent of the number of reasoning steps since the model only learns the induction function. Consequently, the model naturally generalizes to inputs requiring different numbers of reasoning steps. This gives improvements of OOD/length generalization for the cycle task (Figure 4(b)), parity (Figure 5(a)), and addition (Figure 5(b)).
Another interesting aspect is whether the model can use an inductive behavior on new tasks if it was pretrained on prior inductive tasks. Note that the inductive behavior of the inductive scratchpad is only determined by two special tokens. Thus, in principle, models can generate these special tokens and go into the inductive state for other tasks if pretrained on inductive data. We leave the general investigations of pretrained models and the automated learning of more general scratchpads, capitalizing on the measures defined here, to future works.
@@ A.1 Reasoning capabilities of Transformers
Nye et al. [32] put forward the idea of using scratchpads by showing that training Transformers on the intermediate steps in addition to the final answer can improve their performance on different tasks such as executing Python code, addition, and evaluating polynomials. Similarly, in chain-of-though reasoning (CoT) [60] models are shown step-by-step demos of problem-solving (or models generate chains of thought without any examples as in zero-shot CoT [61], among other variants). It has been further shown that using explicit explanations and reducing ambiguity can be proven useful as in the notion of algorithmic prompting [62]. Lanchantin et al. [63] introduced the concept of self-notes, showing that interleaving the intermediate reasoning steps within the question/context rather than after the question can boost the performance of scratchpad/CoT on certain tasks. Goyal et al. [64] introduced pause tokens which act as dummy tokens that provide models with more compute time and processing before the generation of the true next token. On a related note, the iterative prompting method in [65] involves querying LLMs iteratively to optimize question prompts. Further, the Select and Inference framework (SI) [66] utilizes pre-trained LLMs as general processing modules, alternating between selection and inference to generate interpretable reasoning steps leading to the final results. In this work, we used the concept of autoregressive locality to explain why scratchpads are helpful and how to design them. Further, we introduced inductive scratchpads that are suitable for a broad range of algorithmic reasoning tasks and can potentially generalize to more complex samples than what appears in the train distribution.
@@ A.2 Learning measures and lower-bounds for GD
We now discuss the proof techniques for Theorem 1. We prove that the Transformer cannot learn to distinguish between $1$ cycle and $3$ cycles by means of a statistical query argument. We find a group of permutations that preserve the input distribution, take the orbit of the target function under these permutations, and show that a random pair of functions in the orbit are nearly uncorrelated. Then we argue that that means that no function is significantly correlated with a random function in the orbit, so the gradient is uninformative and the net fails to learn anything meaningful. The main complication to this argument is the fact that the input distribution is fairly complicated, in contrast to orbit arguments used in previously discussed works (e.g., for subset parities). The majority of the symbols in the input are fixed and the rest have nontrivial dependencies. However, we get around that by dividing the input into blocks and observing that switching the nonfixed symbols in any two blocks leaves the overall probability distribution unchanged. This argument would largely carry over to other cases with a sufficiently symmetric input distribution and high locality target function.
@@ A.3 On RASP and RASP-L
Also, RASP-L [68] is a variant of RASP that does not contain the full parity function as a short program. Our locality theory predicts that the full parity can be learned by some regular Transformer (in contrast to subset parities), so this also gives a nuance. Incidentally, our reason for expecting that some regular Transformer can learn the parity is that if the Transformer is set to mostly ignore the positional embeddings then in the first layer it will essentially average all of the inputs together. The correct label is a function of this average and it only has $n+1$ possible values, so it seems likely that with the right setup it would be able to memorize the appropriate response to each of them.
@@ B.2 Change of distribution and curriculum learning
We change the training distribution to a uniform mixture of the cycle task distribution for sizes $i=2,\ldots,n$, i.e., each sample comes from the distribution of the cycle task for size $i$ (having $2i$ nodes and edges) with probability $\frac{1}{n-1}$. Therefore, in the mixed distribution setting, the number of reasoning steps varies between $2$ to $n$. As an example, we set $n=7$ and train our model with fresh samples from the mixed distribution. We assess the model’s performance on cycle task samples of sizes $72,\ldots,n=$ 7 (where we include both connected and disconnected cases as in the original definition). The results are shown in Figure 8(a). It can be seen that the cycle task is learned in the order of difficulty (i.e., size). Further, note that when a mixed distribution is not used, weak learning of the cycle task for $n=7$ is not possible up to $100K$ iterations (see also Figure 1(b)), whereas here, weak learning for $n=7$ begins in the first $30K$ iterations. Note that this is not in contrast with our theoretical results since including easy samples reduces the locality. Analogously, in the setting of learning parities, it has been shown that using a biased rather than uniform distribution (which makes the distribution simpler) can make parity targets easier to learn [86, 87].
@@ B.3 Learning parities with scratchpad
Consider $n$ bits $x_{1},\ldots,x_{n}\in\{-1,+1\}$ with a uniform distribution over the Boolean hypercube. As discussed in Section 3.1, learning the parity of any fixed $\frac{n}{2}$ bits is exponentially hard (e.g., an exponential number of iterations in $n$ if the model size is fixed). Note that the locality is $\frac{n}{2}$ in this setting. Here, we particularly focus on learning the parity of the first $\frac{n}{2}$ bits with a scratchpad of cumulative parities as discussed in Section 3.1. In other words, we design a scratchpad given by the sequence
@@
where the autoregressive locality is $2$. With this scratchpad of cumulative parities, we expect the half-parity function to be learned efficiently. We confirm the latter empirically in Figure 10. It can be seen that our decoder-only Transformer can efficiently learn half-parities of growing sizes (e.g., parity of the first 200 bits from 400 bits is the rightmost data point). We note that we have repeated each experiment for $10$ different random seeds and we observed that the number of iterations depends on the the random seed of the experiment. Nevertheless, for all of the seeds, the parities are learned efficiently with a low number of iterations.

### v2

@@ 2.3 Agnostic scratchpads cannot break the globality
Consider training a $T$-regular net with an $s$-scratchpad to learn $P_{X,Y}$ on a constant-sized alphabet by means of the following SGD algorithm. At each timestep, we draw a random sample $(X,Y)$ and compute a value for the net with scratchpad on $X$. Let $\mathcal{S}$ be the resulting scratchpad and for each $i\leq s$ and each $\sigma\neq\mathcal{S}_{i}$, define an alternative scratchpad by setting the first $i-1$ entries of this scratchpad equal to those of $\mathcal{S}$, setting its $i$-th entry to $\sigma$, and using the net to compute the rest of its values. Then, regard the loss associated with setting the $i$th entry of the scratchpad to $\sigma$ as the loss resulting from the associated scratchpad, and use the resulting gradient to carry out one step of SGD. Then if $P_{X}$ is a well-behaved probability distribution, $P_{X,Y}$ is efficiently weakly learnable by a T-regular neural network with a scratchpad if and only if $P_{X,Y}$ has constant globality.
@@
3 Scratchpads to break the globality
@@ 3.1 Educated scratchpad
We now provide a quantitative understanding of how the scratchpad can help with the notion of globality in the autoregressive setting (Definition 3). Assume that we want to learn target $Y\in\mathcal{A}$ from input $X\in\mathcal{A}^{n}$ such that $(X,Y)\sim D$ and $\mathrm{glob}(D)$ is high. If one can design intermediate targets $Y_{1},\ldots,Y_{k}\in\mathcal{A}$ such that $Y_{k}=Y$ and the sequence $(X,Y_{<i})\to Y_{i}$ has low globality according to Definition 3, then one can expect to learn each step of the sequence efficiently and thus the target at the end. In this case, the intermediate targets give the ‘educated scratchpad’ (see Figure 3 for an illustration). We now show how designing low-globality scratchpads can help with learning by focusing on two examples: parity functions and the cycle task.
Figure 3: An illustration showing how scratchpads can break the globality. The target may be efficiently learned if each scratchpad step is of low globality given the previous ones.
@@
Consider learning parity function $y=f(x_{1},\ldots,x_{n})=x_{1}x_{2}\cdots x_{k}$ where $x_{1},\ldots,x_{n}$ are drawn i.i.d. in {±$\{\pm 1\}$ with uniform distribution. For $k\leq\frac{n}{2}$, one can easily check that the globality of this task is $k$ as any $k-1$ coordinates are independent of the output and the histogram of the tokens does not help. Parity functions are known to be hard to learn [12]. More specifically, it has been previously shown that as $k$ increases the parity task becomes harder to learn to the point that parity of degree $k\wedge(n-k)=\omega(1)$ cannot be learned in $poly(n)$ time with standard poly-size neural networks under standard training assumptions [12]. Note that this is consistent with our results, as the globality is non-constant.
@@
The parity task with the cumulative product scratchpad has a globality of $2$.
@@
Consider the cycle task and a scratchpad that learns the depth-first search (DFS) algorithm from the source query node. [^11] For the particular graph structure of the cycle task, DFS is the same as the breadth-first search (BFS). For example, consider the following input corresponding to two cycles $a,x,q$ and $n,y,t$: a>x; n>y; q>a; t>n; y>t; x>q; a?t;. In this case, doing a DFS from node a gives a>x>q>a where the fact that we have returned to the source node a and not seen the destination t indicates that the two nodes are not connected. Therefore, the full scratchpad with the final answer can be designed as a>x>q>a;0. Similarly, if the two nodes were connected the scratchpad would be a>...>t;1. One can easily check that the cycle task becomes low-globality with the DFS scratchpad.
@@
The cycle task with the DFS scratchpad has a globality of $3$.
@@
If one has full knowledge of the target function, one could break the target into sub-targets using an educated scratchpad to keep the globality low and thus learn more efficiently (of course one does not have to learn the target under full target knowledge, but one may still want to let a model learn it to develop useful representations in a broader/meta-learning context). One could in theory push this to learning any target that is poly-time computable by emulating a Turing machine in the steps of the scratchpad to keep the overall globality low. Some works have derived results in that direction, such as [33] for some type of linear autoregressive models, or [12] for more abstract neural nets that emulate any Turing machine with SGD training. However, these are mostly theory-oriented works. In practice, one may be instead interested in devising a more ‘generic’ scratchpad. In particular, a relevant feature in many reasoning tasks is the power of induction. This takes place for instance in the last 2 examples (parity and cycle task) where it appears useful to learn inductive steps when possible.
@@ 3.2 Inductive Scratchpads
As discussed previously, scratchpads can break the global reasoning barrier with appropriate mid-steps. In this part, however, we show that fully educated scratchpads can be sensitive to the number of reasoning steps, translating into poor out-of-distribution (OOD) generalization. As a remedy, we put forward the concept of inductive scratchpad which applies to a variety of reasoning tasks as in previous sections.
@@ 3.2.2 Inductive scratchpad: definition and experimental results
We can use inductive scratchpads to achieve length generalization for the parity and addition tasks. For parities, we insert random spaces between the bits and design an inductive scratchpad based on the position of the bits and then compute the parity iteratively. Using this scratchpad we can train a Transformer on up to $30$ bits and get generalization on up to $50$, or $55$ bits. Results are provided in Figure 5(a). For the addition task, we consider two inductive scratchpad formats. We provide an inductive scratchpad that requires random spaces between the digits in the input and uses the position of the digits to compute the addition digit-by-digit (similar to the parity). With this scratchpad, we can generalize to numbers with 20 digits while training on numbers with up to 10 digits. In the other format, we use random tokens in the input and we compute the addition digit-by-digit by shifting the operands. The latter enables us to generalize from 4 to 26 digits at the cost of having a less natural input format. The results for different seeds are provided in Figure 5(b). See details of these scratchpads in Appendices B.4B.5.121212Our code is also available at https://github.com/aryol/inductive-scratchpad. Also see Appendix A for a comparison with recent methods.
@@ 4 Conclusion
This paper shows that for the learning objective and in contrast to expressivity results, Transformers trained from scratch have a ‘global reasoning barrier’ quantified by the globality degree: if the target requires a global composition of the tokens, Transformers will not efficiently learn. The globality measure has a simpler form and broader applicability range than prior measures as discussed in Appendix A.2, it also has tighter applicability for Transformers. The measure is currently defined for weak learning (inverse-polynomial or constant edge), and a natural next step is to consider stronger learning requirements with notions of ‘globality leap’, e.g., expanding the current work in the direction of [28] but for more general distributions. Investigating the role of curriculum learning is another natural direction and we provide preliminary results here in Appendix B.2.
The globality is also defined in the autoregressive setting, to better quantify when scratchpads can break targets into easier sub-targets. Two negative results are shown for scratchpads: agnostic scratchpads still suffer from the globality barrier, and fully educated scratchpads can have poor OOD generalization. This motivates the introduction of the inductive scratchpad.
The inductive scratchpad can be used for a broad range of reasoning/algorithmic tasks and can easily be integrated into Transformers. The inductive scratchpad is independent of the number of reasoning steps since the model only learns the induction function. Consequently, the model thus better generalizes to inputs requiring different numbers of reasoning steps. This gives improvements of OOD/length generalization for the cycle task (Figure 4(b)), parity (Figure 5(a)), and addition (Figure 5(b)).
Another interesting aspect is whether the model can use an inductive behavior on new tasks if it was pre-trained on prior inductive tasks. Note that the inductive behavior of the inductive scratchpad is only determined by two special tokens. Thus, in principle, models can generate these special tokens and go into the inductive state for other tasks if pre-trained on inductive data. We leave the general investigations of pre-trained models and the automated learning of more general scratchpads, capitalizing on the measures defined here, to future works.
@@ A.1 Reasoning capabilities of Transformers
Nye et al. [32] put forward the idea of using scratchpads by showing that training Transformers on the intermediate steps in addition to the final answer can improve their performance on different tasks such as executing Python code, addition, and evaluating polynomials. Similarly, in chain-of-though reasoning (CoT) [60] models are shown step-by-step demos of problem-solving (or models generate chains of thought without any examples as in zero-shot CoT [61], among other variants). It has been further shown that using explicit explanations and reducing ambiguity can be proven useful as in the notion of algorithmic prompting [62]. Lanchantin et al. [63] introduced the concept of self-notes, showing that interleaving the intermediate reasoning steps within the question/context rather than after the question can boost the performance of scratchpad/CoT on certain tasks. Goyal et al. [64] introduced pause tokens which act as dummy tokens that provide models with more compute time and processing before the generation of the true next token. On a related note, the iterative prompting method in [65] involves querying LLMs iteratively to optimize question prompts. Further, the Select and Inference framework (SI) [66] utilizes pre-trained LLMs as general processing modules, alternating between selection and inference to generate interpretable reasoning steps leading to the final results. In this work, we used the concept of autoregressive globality to explain why scratchpads are helpful and how to design them. Further, we introduced inductive scratchpads that are suitable for a broad range of algorithmic reasoning tasks and can potentially generalize to more complex samples than what appears in the train distribution.
@@ A.2 Learning measures and lower-bounds for GD
We now discuss the proof techniques for Theorem 1. We prove that the Transformer cannot learn to distinguish between $1$ cycle and $3$ cycles by means of a statistical query argument. We find a group of permutations that preserve the input distribution, take the orbit of the target function under these permutations, and show that a random pair of functions in the orbit are nearly uncorrelated. Then we argue that that means that no function is significantly correlated with a random function in the orbit, so the gradient is uninformative and the net fails to learn anything meaningful. The main complication to this argument is the fact that the input distribution is fairly complicated, in contrast to orbit arguments used in previously discussed works (e.g., for subset parities). The majority of the symbols in the input are fixed and the rest have nontrivial dependencies. However, we get around that by dividing the input into blocks and observing that switching the nonfixed symbols in any two blocks leaves the overall probability distribution unchanged. This argument would largely carry over to other cases with a sufficiently symmetric input distribution and high globality target function.
@@ A.3 On RASP and RASP-L
Also, RASP-L [68] is a variant of RASP that does not contain the full parity function as a short program. Our globality theory predicts that the full parity can be learned by some regular Transformer (in contrast to subset parities), so this also gives a nuance. Incidentally, our reason for expecting that some regular Transformer can learn the parity is that if the Transformer is set to mostly ignore the positional embeddings then in the first layer it will essentially average all of the inputs together. The correct label is a function of this average and it only has $n+1$ possible values, so it seems likely that with the right setup it would be able to memorize the appropriate response to each of them.
@@ B.2 Change of distribution and curriculum learning
We change the training distribution to a uniform mixture of the cycle task distribution for sizes $i=2,\ldots,n$, i.e., each sample comes from the distribution of the cycle task for size $i$ (having $2i$ nodes and edges) with probability $\frac{1}{n-1}$. Therefore, in the mixed distribution setting, the number of reasoning steps varies between $2$ to $n$. As an example, we set $n=7$ and train our model with fresh samples from the mixed distribution. We assess the model’s performance on cycle task samples of sizes $72,\ldots,n=$ 7 (where we include both connected and disconnected cases as in the original definition). The results are shown in Figure 8(a). It can be seen that the cycle task is learned in the order of difficulty (i.e., size). Further, note that when a mixed distribution is not used, weak learning of the cycle task for $n=7$ is not possible up to $100K$ iterations (see also Figure 1(b)), whereas here, weak learning for $n=7$ begins in the first $30K$ iterations. Note that this is not in contrast with our theoretical results since including easy samples reduces the globality. Analogously, in the setting of learning parities, it has been shown that using a biased rather than uniform distribution (which makes the distribution simpler) can make parity targets easier to learn [86, 87].
@@ B.3 Learning parities with scratchpad
Consider $n$ bits $x_{1},\ldots,x_{n}\in\{-1,+1\}$ with a uniform distribution over the Boolean hypercube. As discussed in Section 3.1, learning the parity of any fixed $\frac{n}{2}$ bits is exponentially hard (e.g., an exponential number of iterations in $n$ if the model size is fixed). Note that the globality is $\frac{n}{2}$ in this setting. Here, we particularly focus on learning the parity of the first $\frac{n}{2}$ bits with a scratchpad of cumulative parities as discussed in Section 3.1. In other words, we design a scratchpad given by the sequence
@@
where the autoregressive globality is $2$. With this scratchpad of cumulative parities, we expect the half-parity function to be learned efficiently. We confirm the latter empirically in Figure 10. It can be seen that our decoder-only Transformer can efficiently learn half-parities of growing sizes (e.g., parity of the first 200 bits from 400 bits is the rightmost data point). We note that we have repeated each experiment for $10$ different random seeds and we observed that the number of iterations depends on the the random seed of the experiment. Nevertheless, for all of the seeds, the parities are learned efficiently with a low number of iterations.